Find an equation that has the solutions: x= -4 and 7
select 1 answer below:
a.x^2 - 3x - 28 = 0
b.x^2 + 3x = 28
c.x^x - 11x = 28
d.x^2 + 3x + 28 = 0

Answer by checkley71(8403)

A) IS THE ONLY ONE WITH THE PROPER TERMS.
X^2-3X-28=0
(X+4)(X-7)=0
X+4=0
X=-4 ANSWER.
X-7=0
X=+7 ANSWER.